### v2

@@ Abstract
With the large-scale integration of Internet of Things (IoT) into enterprise information management systems, organizations are pursuing digital transformation that hinges on real-time data insights—and yet face escalating security and governance risks. Detecting and responding to threats at scale without impairing system efficiency has therefore become a critical information-management and decision-support challenge for today’s executives. This paper develops a distributed, gain-based anomaly-detection framework tailored to IoT-enabled enterprise systems, underpinned by an optimized sensor-subset partitioning strategy. Starting from the perspective of set partitioning strategies, this study analyzes the key factor that contributes to the performance differences between distributed and centralized algorithms. By examining the gain mutual influence of sensor subsets, an optimal set partitioning strategy is designed to minimize inter-subset mutual influence while enhancing intra-subset correlation. To further reduce the computational cost of gain updates, a suboptimal partitioning strategy based on Grassmann distance is proposed, improving the efficiency of selecting suspicious sensors. Theoretical analysis demonstrates that this approach effectively reduces the computational cost of gain updates while maintaining detection performance. Finally, simulation results validate the effectiveness of the proposed method in enhancing attack detection performance.
@@ 1 Introduction
With the rapid development of information technology, automation and intelligent systems are profoundly reshaping the way modern society operates. As a key supporting technology for this change, sensor networks have been widely used in various fields, including smart cities, the Internet of Things, and industrial manufacturing (Heo et al., 2009; Habibzadeh et al., 2017; Hatami et al., 2022; Wang et al., 2020). By enabling comprehensive environmental monitoring, real-time data acquisition, and autonomous control, these networks significantly enhance the intelligence and decision-making capabilities of enterprise information management systems (Wei et al., 2022; Tao et al., 2022; Ju et al., 2023).
However, the proliferation of sensor deployments introduces complex security and governance challenges, as organizational decision makers must balance efficient system operation with the timely detection and response to emerging threats(Xie et al., 2018). Especially in open and dynamic network environments, adversaries may interfere with the system’s state estimation by means of data tampering and information deception, resulting in system malfunctions or even serious security incidents, bringing significant risks to infrastructure and key businesses (Ge et al., 2019; Pang et al., 2021; Smith et al., 2021). Therefore, in the context of large-scale distributed sensor networks, the development of an efficient and accurate security defense mechanism has emerged as a crucial research direction to ensure system stability and reliability.
@@
Attack detection has long been a core area of research in cybersecurity, encompassing critical domains such as intrusion detection, anomaly detection, and fault diagnosis (Wang et al., 2025; Suo et al., 2024a; Li et al., 2023, 2024; Zhang et al., 2021; Zhao et al., 2023; Balta et al., 2023). Existing methods primarily fall into two categories: rule-based and learning-based approaches. Rule-based approaches rely on predefined security policies and pattern matching techniques, such as anomaly threshold settings and signature detection. For example, Wang et al. (2025) extracted topological relationships measured by individual nodes and used reconstruction residuals to pinpoint the location of injection attacks. Suo et al. (2024a) explored the Grassmann distance of opinion evolution direction and leveraged its latent characteristics to iteratively filter out malicious agents. Li et al. (2023) constructed a dynamic relationship between raw data and decrypted results to enable rapid detection of malicious attacks. In the realm of manufacturing system security, Li et al. (2024) developed detection rules based on processing procedures and key parameters, identifying network attacks through rule matching. However, rule-based approaches struggle to adapt to novel and complex attacks. In contrast, learning-based approaches leverage data-driven techniques such as statistical analysis, machine learning, and deep learning to detect and recognize attack patterns, offering superior generalization and adaptability (Zhang et al., 2021). For instance, Zhao et al. (2023) designed a data-driven attack detector based on subspace identification, detecting attacks by computing the system’s stable kernel representation. Balta et al. (2023) proposed a detection framework based on digital twins, identifying attacks by analyzing controlled transient behaviors. To address traffic anomaly detection, Zhang et al. (2022) designed a model consisting of two adversarial sub-networks to learn the data distribution of normal traffic. And based on this, an anomaly detection method based on high anomaly suppression is proposed. Nevertheless, existing research indicates that these methods still face significant challenges in model training and data dependency (Li et al., 2024; Zhang et al., 2021). Thus, improving adaptability and computational efficiency while ensuring detection accuracy remains a critical research direction in attack detection.
@@
For large-scale networks, a sufficient condition to ensure system security is that each sensor can directly exclude potential attacked sensors from the neighbor set. For instance, the ADS algorithm proposed in literature Suo et al. (2024c) aims to achieve this objective. However, the algorithm is centralized for each node, which means that its efficiency may decrease as the neighborhood cardinality increases. Given that a distributed approach can enhance the efficiency of a centralized algorithm, this paper explores the feasibility of extending the ADS algorithm into a distributed form. Existing studies indicate that distributed algorithms generally lag behind centralized ones in performance, with partitioning strategies considered one of the key factors influencing this gap (Mirzasoleiman et al., 2016).
Existing studies have shown that partitioning large-scale sets into subsets based on feature characteristics is an effective way to address the curse of dimensionality (Wang et al., 2023). The key challenge lies in identifying and selecting informative features. To more effectively design partitioning strategies, this paper will extract the inherent characteristics of each sensor. Based on this, the paper aims to optimize the design of partitioning strategies to balance the combined effect of inter-subset benefits and intra-subset correlations. Through optimized partitioning strategy design, the goal of this paper is to narrow the performance gap between distributed and centralized ADS algorithms, achieving a trade-off between efficiency and performance in the suspicious sensor selection task.
The main contributions of this paper are summarized as follows: The concept of mutual influence is introduced to quantify the marginal benefit interference between different sensor subsets. Based on the observable space characteristics of sensors, this metric effectively guides the set partitioning process, aiming to minimize inter-subset mutual influence. To address the imbalance in sensor gain updates caused by differences in observable spaces within subsets, a suboptimal set partitioning strategy based on Grassmann distance is proposed. This strategy maximizes the correlation among elements within each subset while minimizing mutual influence between subsets, providing strong support for extending the ADS algorithm to its distributed form (D-ADS). Under the proposed set partitioning strategy, gain updates are required only within the selected subset, significantly reducing computational costs. Theoretical results demonstrate that the error in the probability distribution vector remains bounded, ensuring the detection performance of suspicious sensor selection while improving algorithm efficiency. Finally, extensive simulation results under various scenarios validate the effectiveness of the proposed algorithm.
@@
This paper is organized as follows: Section II presents the system model and problem formulation, including the description of the sensor network, attacker model, and the suspicious sensor selection problem in dynamic environments. Section III focuses on the analysis of the key factors affecting the performance of distributed algorithms, particularly the mutual influence of benefits between subsets and the design of the set partitioning strategy based on Grassmann distance. Section IV evaluates the performance of the proposed strategies through numerical simulations. Finally, Section V concludes the paper by summarizing the main contributions, and suggesting potential directions for future research.
@@ 2.1 System Model
where $y_{i}\left(k\right)\in\mathbb{R}^{m}$ and $\nu_{i}\left(k\right)\in\mathbb{R}^{m}$ represent the measurement and measurement noise of sensor $i$, respectively. The matrix $A\in\mathbb{R}^{n\times n}$ in equation (1) and the matrix $C_{i}\in\mathbb{R}^{m\times n}$ in equation (2) are both real matrices with appropriate dimensions. Similarly, $\nu_{i}\left(k\right)$ follows a Gaussian distribution with zero mean and a positive definite covariance matrix $R_{i}$, i.e., $\nu_{i}\left(k\right)\sim\mathcal{N}\left(0,R_{i}\right)$. In this section, both the process noise $||\omega\left(k\right)||$ and the measurement noise $||\nu_{i}\left(k\right)||$ are upper-bounded by certain positive scalars.
@@ 2.2 Attacker Model
For the attacker’s strategy, the following assumptions are made:
  • At any moment $k$, the number of compromised sensors $q_{i}$ in the neighborhood of sensor $i$ does not exceed half the size of the neighborhood, i.e., $q_{i}\leq\lfloor|\mathcal{N}_{i}|/2\rfloor$Yang and Lv (2021); Lu and Yang (2023a)
@@
2.3 Suspicious Sensor Selection Problem
Lemma 1 in literature Suo et al. (2024c) states that selecting the suspicious sensor set from the set $\mathcal{N}_{i}$ is an NP-hard problem and that the objective function can be transformed into a submodular function as shown in equation (6)
@@ 2.3 Suspicious Sensor Selection Problem
where $\varLambda_{i,k}$ is the augmented error matrix, which is obtained by summarizing the estimated error $||\hat{x}_{i}(k)-\hat{x}_{ij}^{a}(k)||$, i.e., $\varLambda_{i,k}=\sum^{|\mathcal{N}_{i}|}_{j=1}{\left(\theta_{|\mathcal{N}_{i}|}^{j}\otimes||\hat{x}_{i}\left(k\right)-\hat{x}_{ij}^{a}\left(k\right)||\right)}$, and $[\mu_{ij}(k)]_{j\in\mathcal{N}_{i}}$ denotes the augmented matrix of $\mu_{ij}(k)$, where $\mu_{ij}(k)=1$ for $j\in\mathcal{A}_{i,k}$ and $\mu_{ij}(k)=0$ for $j\in\mathcal{N}_{i}\backslash\mathcal{A}_{i,k}$. Consequently, the suspicious sensor selection problem is reformulated as a solvable submodular maximization problem under a cardinality constraint, as presented in Problem 1.
@@
Suo et al. (2024c) The problem of selecting suspicious sensors essentially involves choosing no more than $q_{i}$ sensors to maximize the objective function (6), i.e.,
$\max_{\mathcal{A}_{i,k}\subseteq\mathcal{N}_{i}}f_{k}\left(\mathcal{A}_{i,k}\right),\leavevmode\nobreak\ \leavevmode\nobreak\ \text{s.t.}\left|\mathcal{A}_{i,k}\right|\leq q_{i},$ (7)
@@ 3 Main Results
It is shown that distributed algorithms have obvious advantages over centralized algorithms in terms of efficiency, but from the perspective of performance, distributed algorithms generally have a certain gap with centralized algorithms. Therefore, this section focuses on the following issues: analyzing the key factors that affect the performance difference between D-ADS and ADS algorithm, and designing effective partitioning strategies to minimize the performance gap.
@@ 3.1 The mutual influence of benefits between subsets
In this section, the neighbor set of each sensor is divided into several subsets, and the ADS algorithm in the literature Suo et al. (2024c) is deployed on each subset. At this time, the ADS algorithm becomes a distributed case, that is, the D-ADS algorithm.
Specifically, the neighbor set $\mathcal{N}_{i}$ of sensor $i$ is divided into $m_{i}$ subsets, expressed as $\mathcal{N}_{i}=\mathcal{N}_{i,1}\cup\mathcal{N}_{i,2}\cup\cdots\cup\mathcal{N}_{i,m_{i}}$. It should be noted that the parameter $m_{i}$ is predetermined and has nothing to do with $|\mathcal{N}_{i}|$.
The method of approximate average partitioning is adopted to ensure that the cardinality of each subset is approximately equal [^1] At this point, the balance between the number of attacked sensors in each subset and the attack intensity can be obtained intuitively. And the results from literature Suo et al. (2024c) indicate that the D-ADS algorithm guarantees a performance lower bound for each subset consistent with that of the ADS algorithm. The proof process is omitted here. Please refer to the Lemma 2 in APPENDIX A and B.. The cardinality $|\mathcal{N}_{i,g_{i}}|$ of the $g_{i}$-th subset can be expressed as:
@@
where $g_{i}=1:m_{i}$, and the indicator function $\mathbb{I}(\cdot)$ is defined as follows: when $g_{i}$ is less than or equal to the remainder of $|\mathcal{N}_{i}|$ divided by $m_{i}$, $\mathbb{I}(\cdot)$ is $1$, otherwise $\mathbb{I}(\cdot)$ is 00.
However, to ensure this performance lower bound, the gain values of all sensors need to be updated after each selection, and existing literature shows that this will bring huge computational complexity Mirzasoleiman et al. (2016). In fact, at the $l$-th selection at moment $k$, only the sensor $j_{g_{i}^{s},select}^{l_{g_{i}^{s}}}$ in the $g_{i}^{s}$-th subset is selected. To reduce the amount of calculation, this section starts with the gain update method and explores the feasibility of only updating the gains of the remaining elements of the subset where the selected sensor is located.
@@
By calculating the indicator function for each row of $Q_{j,o}$, a column vector $[I_{j,\ell}]_{\ell=1:n}\in\mathbb{R}^{n}$ can be obtained. Then, the mutual influence is defined in Definition 2:
@@
(Mutual Influence) For any two sensors $j_{g_{i}}$ and $j_{q_{i}}$ in any neighbor sensor subsets $\mathcal{N}_{i,k,g_{i}}$ and $\mathcal{N}_{i,k,q_{i}}$ of sensor $i$, the mutual influence of gains between them is defined as the number of dimensions $[I_{j_{g_{i}},\ell}]_{\ell=1:n}$ and $[I_{j_{q_{i}},\ell}]_{\ell=1:n}$ that are not 00 simultaneously in all dimensions $\ell=1:n$, denoted as
@@
During the selection of a suspicious sensor, assume that the sensor $j_{g_{i}}$ from the $g_{i}$-th neighbor subset of sensor $i$, i.e., $j_{g_{i}}\in\mathcal{N}_{i,k,g_{i}}$, is selected. According to equation (9), the $Q_{j,o}$ of each sensor $j$ can be transformed into an indicator function vector. Therefore, it is only necessary to ensure that mutual influence between the elements in other subsets $q_{i}$ and the remaining elements of subset $g_{i}$ is minimized, that is,
@@
At this point, an optimal set partitioning algorithm based on the observable discrimination matrix can be proposed, which can be calculated offline.
The observable part of each sensor is extracted based on the observable discrimination matrix, so the mutual influence of gain only occurs between sensors with intersections in the observable space. Therefore, based on the optimal set partitioning strategy, after each suspicious sensor is selected, only the gains of the sensors in the subset need to be updated. However, the observable spaces of the sensors in the same subset may not be exactly the same. This means that after each sensor selection, the number of dimensions that need to be updated in the estimated error vector of each sensor in the subset will be different, resulting in an inaccurate allocation ratio vector $p_{p,k}^{(l_{p})}$.
3.2 The Suboptimal set partitioning strategy based on Grassmann distance
Based on the previous analysis, this paper considers exploring a suboptimal partitioning strategy to strike a balance between the mutual influence between subsets and the correlation within subsets.

### v3

@@ Abstract
[…]
@@ 1 Introduction
1.1 Background and Related works
The rapid development of the IoT is driving digital transformation across key sectors such as energy, power, healthcare, and smart manufacturing. Billions of heterogeneous devices continuously collect and transmit massive volumes of data through sensors and network interfaces, making IoT a fundamental enabler of intelligent control and automation (Zhang and Tao, 2020).
However, large-scale connectivity, device heterogeneity, and long life cycles expose unique attack surfaces across the device–edge–network–cloud continuum, posing unprecedented security challenges. Resource constraints and “headless deployments” complicate patch management, while issues such as default credentials, insecure firmware updates, and expired certificates remain widespread (Yousefnezhad et al., 2020). Lightweight protocols (e.g., MQTT, CoAP), if lacking encryption and key rotation, are highly vulnerable to man-in-the-middle and replay attacks (Mathews and Gondkar, 2019). At the Information Technology(IT)/Operational Technology(OT) convergence boundary, IoT systems must balance real-time performance with reliability, where security failures can escalate directly into risks at the physical layer (Pascoe, 2023). Ensuring system reliability and data trustworthiness in such large-scale, heterogeneous environments has therefore become a central challenge for the widespread adoption of IoT applications (Alsalem and Amin, 2023).
At the framework and methodology level, several widely adopted cybersecurity frameworks provide foundational guidance for IoT security. NIST CSF defines an organizational risk management cycle through five core functions, includes Identify, Protect, Detect, Respond, and Recover (Pascoe, 2023; Aljumaiah et al., 2025). IEC 62443, tailored for industrial control and OT systems, highlights layered and defense-in-depth strategies (Leander et al., 2019). ISO/IEC 27001 and 30141 offer general guidance on information security management and IoT reference architectures (Humphreys, 2016; Sugiharto and Kaburuan, 2023), while ENISA’s threat landscape reports regularly update major IoT threats and mitigation trends across Europe and beyond (European Union Agency for Network and Information Security (2019), ENISA). Despite their value, these frameworks largely operate at a strategic and governance level, with limited emphasis on dynamic attack detection and response mechanisms required in practice.
Beyond the general governance frameworks proposed by standardization bodies, the academic community has also explored security frameworks tailored to IoT scenarios. Halgamuge and Niyato (2025) introduced an adaptive edge security framework capable of dynamically generating security policies to address complex and evolving threats. To tackle security issues arising from IoT’s heterogeneous architectures, Masud et al. (2025) proposed a hybrid moving target defense (MTD)-based security level analysis method for assessing network states. Pavithran et al. (2020) identified key elements for building secure blockchain architectures in IoT, contributing to improved system robustness. Addressing external intrusion risks, Qaddos et al. (2024) designed a deep learning architecture capable of capturing complex features, offering potential for safeguarding IoT against security threats. Overall, these studies have achieved notable progress in authentication, protection, and anomaly detection, and to some extent help bridge the gap between high-level standardized frameworks and their practical implementation.
Sensor networks form a core component enabling collaboration among heterogeneous IoT devices. However, as the scale of IoT systems continues to expand, sensor networks themselves are growing at an exponential rate. This trend introduces complex security and governance challenges, as decision-makers must balance efficient system operation with the timely detection and response to emerging threats (Xie et al., 2018). Adversaries may disrupt state estimation through data tampering or information deception, leading to system malfunctions and even severe security incidents that pose significant risks to critical infrastructure and essential services (Ge et al., 2019; Pang et al., 2021; Smith et al., 2021). Although end-to-end security mechanisms, such as schemes based on pre-shared keys or digital certificates, have been deployed in some sensor networks, the compromise of keys or certificates can still result in data tampering and integrity breaches (Kwon et al., 2013; Mouha and Mouha, 2021). Moreover, despite substantial progress in IoT security research, studies on large-scale distributed sensor networks remain limited, particularly with respect to addressing complex integrity attacks that are both dynamic and stealthy. Developing efficient and precise defensive mechanisms has therefore become a critical research direction for ensuring the stability and reliability of such systems.
@@ 1.1 Background and Related works
Attack detection has long been a core area of research in cybersecurity, encompassing critical domains such as intrusion detection, anomaly detection, and fault diagnosis (Wang et al., 2025; Suo et al., 2024a; Li et al., 2023, 2024; Zhang et al., 2021; Zhao et al., 2023; Balta et al., 2023). Existing methods primarily fall into two categories: rule-based and learning-based approaches. Rule-based approaches rely on predefined security policies and pattern matching techniques, such as anomaly threshold settings and signature detection. For example, Wang et al. (2025) extracted topological relationships measured by individual nodes and used reconstruction residuals to pinpoint the location of injection attacks. For the situation where malicious agents are in a dominant position, Suo et al. (2024a) explored using latent features to iteratively filter out malicious agents. Li et al. (2023) constructed a dynamic relationship between raw data and decrypted results to enable rapid detection of malicious attacks. In the realm of manufacturing system security, Li et al. (2024) developed detection rules based on processing procedures and key parameters, identifying network attacks through rule matching. However, rule-based approaches struggle to adapt to novel and complex attacks. In contrast, learning-based approaches leverage data-driven techniques such as statistical analysis, machine learning, and deep learning to detect and recognize attack patterns, offering superior generalization and adaptability (Zhang et al., 2021). For instance, Zhao et al. (2023) designed a data-driven attack detector based on subspace identification, detecting attacks by computing the system’s stable kernel representation. Balta et al. (2023) proposed a detection framework based on digital twins, identifying attacks by analyzing controlled transient behaviors. To address traffic anomaly detection, Zhang et al. (2022) designed a model consisting of two adversarial sub-networks to learn the data distribution of normal traffic. And based on this, an anomaly detection method based on high anomaly suppression is proposed. Nevertheless, existing research indicates that these methods still face significant challenges in model training and data dependency (Li et al., 2024; Zhang et al., 2021). Thus, improving adaptability and computational efficiency while ensuring detection accuracy remains a critical research direction in attack detection.
@@
1.2 Motivation and Contributions
In large-scale sensor networks, each sensor is typically connected to a substantial number of neighbors. In such scenarios, if every sensor has to sequentially identify potential malicious information from its entire neighborhood, the computational overhead becomes prohibitively high and the detection efficiency drops significantly (Suo et al., 2024c). Therefore, achieving both system security and computational efficiency within local neighborhoods remains a critical and unsolved problem.
Recent studies have shown that partitioning large-scale datasets and performing parallel screening across subsets can help mitigate the curse of dimensionality and significantly improve efficiency (Wang et al., 2023). Existing clustering and partitioning techniques, however, typically rely on the process of first collecting data and then grouping nodes based on correlation or similarity features (Xia et al., 2021; Mirzasoleiman et al., 2016). In practice, the collected data are often affected by communication noise and external interference, resulting in unstable classification and suboptimal partitioning outcomes (Han et al., 2022). In particular, under malicious attacks, existing classification algorithms are primarily designed to detect the attacks themselves rather than to pre-classify datasets that already contain adversarial information (Ding et al., 2022; Thakkar and Lohiya, 2023). Motivated by these limitations, this paper seeks to explore more inherent and fundamental data features to enable stable and effective grouping in the offline stage.
At the same time, although distributed methods are inherently more suitable for large-scale scenarios in terms of efficiency, they still suffer from significant performance drawbacks compared to centralized approaches (Mirzasoleiman et al., 2016). One key reason is the lack of a systematic understanding of the relationship between partition strategies and node features. To address this issue, this paper further focuses on optimizing partition strategies, with the goal of narrowing the performance gap between centralized and distributed methods after partitioning, thereby achieving a better balance between efficiency and performance in malicious information selection tasks. The main contributions of this paper are summarized as follows:
  1. 1.
    This paper proposes a “mutual influence” metric to quantify the marginal interference among different sensor subsets. This metric provides a quantitative basis for set partitioning, enabling the partition to maximize the intra-subset correlation while minimizing inter-subset interference, thereby revealing the fundamental reason of the performance gap between distributed and centralized methods.
  2. 2.
    This paper designs a general set partitioning strategy that characterizes sensors by their observable spaces and incorporates the Grassmann distance to achieve robust grouping. This strategy demonstrates strong transferability, enabling the extension of existing neighbor selection–based detection algorithms into a distributed framework. Compared with thenon-partitioned counterpart, the distributed form aligns more naturally with parallel computing architectures, thereby providing an efficient and practical solution for attack detection in large-scale, resource-constrained IoT scenarios.
  3. 3.
    In view of the suitability of submodular optimization theory (SOT) for neighbor-selection problems, this paper extends the SOT-based attack detection scheduling (ADS) algorithm (Suo et al., 2024c) into a two-stage Distributed-ADS (D-ADS) algorithm by incorporating the proposed set partitioning strategy. Both theoretical analysis and simulation results consistently demonstrate that the set partitioning strategy substantially reduces computational cost while maintaining the performance gap between distributed and centralized attack detection within a tolerable range.
Therefore, the core contribution of this study lies in the proposed set partitioning strategy, which is integrated into a theoretically guaranteed Attack Detection Scheduling (ADS) algorithm. Building upon this foundation, we further extend the ADS algorithm to design a more cost-effective and efficient Distributed-ADS (D-ADS) algorithm. The relationship between ADS and D-ADS is illustrated in Figure 1.
Figure 1: Relationship diagram between ADS and D-ADS in the malicious information selection task
@@ 1.2 Motivation and Contributions
This paper is organized as follows: Section II presents the system model and problem formulation, including the description of the sensor network, attacker model, and the malicious information selection problem in dynamic environments. Section III focuses on the analysis of the key factors affecting the performance of distributed algorithms, particularly the mutual influence of benefits between subsets and the design of the set partitioning strategy based on Grassmann distance. Section IV evaluates the performance of the proposed strategies through numerical simulations. Finally, Section V concludes the paper by summarizing the main contributions, and suggesting potential directions for future research.
@@ 2.1 System Model
where $y_{i}\left(k\right)\in\mathbb{R}^{m}$ and $\nu_{i}\left(k\right)\in\mathbb{R}^{m}$ represent the measurement and measurement noise of sensor $i$, respectively. The matrix $A\in\mathbb{R}^{n\times n}$ in equation (1) and the matrix $C_{i}\in\mathbb{R}^{m\times n}$ in equation (2) are both real matrices with appropriate dimensions. Similarly, $\nu_{i}\left(k\right)$ follows a Gaussian distribution with zero mean and a positive definite covariance matrix $R_{i}$, i.e., $\nu_{i}\left(k\right)\sim\mathcal{N}\left(0,R_{i}\right)$. In this section, both the process noise $\omega\left(k\right)||$ and the measurement noise $\nu_{i}\left(k\right)||$ are upper-bounded by certain positive scalars.
@@ 2.2 Attacker Model
For the attacker’s strategy, the following assumptions are satisfied
  • At any moment $k$, the number of compromised sensors $q_{i}$ in the neighborhood of sensor $i$ does not exceed half the size of the neighborhood, i.e., $q_{i}\leq\lfloor|\mathcal{N}_{i}|/2\rfloor$(Yang and Lv, 2021; Lu and Yang, 2023a)
@@
2.3 Malicious Information Selection Problem
Existing literature shows that selecting a set of malicious information from the set $\mathcal{N}_{i}$ is an NP-hard problem, and the objective function can be converted into a submodular function (Suo et al., 2024c)
@@ 2.3 Malicious Information Selection Problem
[…]
@@
(Suo et al., 2024c) The problem of selecting malicious information essentially involves selecting no more than $q_{i}$ sensors to maximize the objective function (6), i.e.,
$\max_{\mathcal{A}_{i,\,k}\subseteq\mathcal{N}_{i}}f_{k}\left(\mathcal{A}_{i,\,k}\right),\penalty 10000\ \penalty 10000\ \text{s.t.}\left|\mathcal{A}_{i,\,k}\right|\leq q_{i},$ (7)
@@ 3.1 The mutual influence of benefits between subsets
In this section, the neighbor set of each sensor is divided into several subsets. Specifically, the neighbor set $\mathcal{N}_{i}$ of sensor $i$ is divided into $m_{i}$ subsets, expressed as $\mathcal{N}_{i}=\mathcal{N}_{i,\,1}\cup\mathcal{N}_{i,\,2}\cup\cdots\cup\mathcal{N}_{i,\,m_{i}}$. It should be noted that the parameter $m_{i}$ is predetermined and has nothing to do with $|\mathcal{N}_{i}|$.
The method of approximate average partitioning is adopted to ensure that the cardinality of each subset is approximately equal [^1] At this point, the balance between the number of attacked sensors in each subset and the attack intensity can be obtained intuitively. Please refer to the Lemma 2 in APPENDIX A.. The cardinality $|\mathcal{N}_{i,\,g_{i}}|$ of the $g_{i}$-th subset can be expressed as:
@@
where $g_{i}=1:m_{i}$, and the indicator function $\mathbb{I}(\cdot)$ is defined as follows: when $g_{i}$ is less than or equal to the remainder of $|\mathcal{N}_{i}|$ divided by $m_{i}$, $\mathbb{I}(\cdot)$ is $1$, otherwise $\mathbb{I}(\cdot)$ is 0.
However, to ensure this performance lower bound, the gain values of all sensors need to be updated after each selection, and existing literature shows that this will bring huge computational complexity (Mirzasoleiman et al., 2016). In fact, at the $l$-th selection at moment $k$, only the sensor $j_{g_{i}^{s},\,select}^{l_{g_{i}^{s}}}$ in the $g_{i}^{s}$-th subset is selected. To reduce the amount of calculation, this section starts with the gain update method and explores the feasibility of only updating the gains of the remaining elements of the subset where the selected sensor is located.
@@
By calculating the indicator function for each row of $Q_{j,\,o}$, a column vector $I_{j,\,\ell}]_{\ell=1:n}\in\mathbb{R}^{n}$ can be obtained. Then, the mutual influence is defined in Definition 2:
@@
(Mutual Influence) For any two sensors $j_{g_{i}}$ and $j_{q_{i}}$ in any neighbor sensor subsets $\mathcal{N}_{i,\,k,\,g_{i}}$ and $\mathcal{N}_{i,\,k,\,q_{i}}$ of sensor $i$, the mutual influence of gains between them is defined as the number of dimensions $I_{j_{g_{i}},\,\ell}]_{\ell=1:n}$ and $I_{j_{q_{i}},\,\ell}]_{\ell=1:n}$ that are not 0 simultaneously in all dimensions $\ell=1:n$, denoted as
@@
During the selection of a malicious information, assume that the sensor $j_{g_{i}}$ from the $g_{i}$-th neighbor subset of sensor $i$, i.e., $j_{g_{i}}\in\mathcal{N}_{i,\,k,\,g_{i}}$, is selected. According to equation (9), the $Q_{j,\,o}$ of each sensor $j$ can be transformed into an indicator function vector. Therefore, it is only necessary to ensure that mutual influence between the elements in other subsets $q_{i}$ and the remaining elements of subset $g_{i}$ is minimized, that is,
@@
Based on the observability discrimination matrix, the observable part of each sensor is extracted. As a result, the mutual influence of gains only occurs between sensors whose observable spaces overlap. Therefore, as long as all sensors with overlapping observable spaces are grouped into the same subset, after each malicious information selection, only the gains of the sensors within that subset need to be updated. However, the observable spaces of sensors within the same subset may not be exactly identical. This means that, after each sensor selection, the number of dimensions in the estimation error vector that need to be updated differs across sensors in the subset, which in turn leads to inaccuracies in the allocation ratio vector.
3.2 The set partitioning strategy based on Grassmann distance
Based on the previous analysis, this paper considers exploring the partitioning strategy to strike a balance between the mutual influence between subsets and the correlation within subsets.